A 10.67-cm-long, 216 g rod is pivoted at one end. A 19 g ball of clay is stuck on the other end. What is the period if the rod and clay swing as a pendulum?
Answer #1
Given that
The length of the rod (L) = 10.67cm
                                      = 0.1067m
The mass of the rod (M) = 0.216kg
The mass of the clay (m) = 19g
                                     = 0.019kg
Accceleration due to gravity (g) =9.8m/s2
We know that the time period of the rod and clay using as apendulum
                          T = 2π√I/mgh
Now the moment of inertia is given by
                          I = (1/3)ML2 +mL2
The height reached by the rod and clay swing is given by
                         h = xcm = mx1 + Mx2/ (m+M)
Here the distnace (x1) =L = 0.1067m and x2 = (L/2)= 0.05335m
Now substitute all the above values to get the requiredanswer